OpenStudy (anonymous):

given cos theta =(*radical*15)/4 and csc thta >0, find sin and tan theta

OpenStudy (amistre64):

sin^2 + cos^2 = 1 ; is the pythag thrm for trig functions

OpenStudy (amistre64):

or one of them that is :)

OpenStudy (amistre64):

\[(\frac{\sqrt{15}}{4})^2+(\frac{?}{?})^2=1\] \[\frac{15}{16}+(\frac{?}{?})^2=1\] \[(\frac{?}{?})^2=1-\frac{15}{16}\] \[(\frac{?}{?})=\sqrt{\frac{1}{16}}\] \[sin(t)=1/4\]

OpenStudy (amistre64):

csc(t) is just sin flipped over; or 4/1

OpenStudy (amistre64):

tan = sin/cos 1/4 --------- = 1/sqrt(15) = sqrt(15)/15 sqrt(15)/4